OpenStudy (anonymous):

2(x-6)/3=x+3 with steps please

OpenStudy (anonymous):

distributing the 2 you get... (2x -12)/3=x+6

OpenStudy (anonymous):

dividing each piece on the left hand side by 3 you get: (2/3)x-4=x+6

OpenStudy (anonymous):

adding the 4 to both sides you get... (2/3)x = x +10

OpenStudy (anonymous):

Subtracting x from both sides you get... (-1/3)x = 10

OpenStudy (anonymous):

odk 

OpenStudy (anonymous):

multiplying both sides by -3/1 ( the reciprocal) gives.. x=-30

OpenStudy (anonymous):

lifesaver.

OpenStudy (anonymous):

glad to help!!